OpenStudy (anonymous):

verify that this is an identity. tan^2 theta +1/sec theta = sec theta

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{tan^2\theta +1}{sec\theta}\\~\\ = \dfrac{sec^2\theta}{sec\theta}\\~\\ = sec\theta}\end{align}\)

OpenStudy (mathmath333):

remember \(\large\tt \begin{align} \color{blue}{tan^2\theta+1=sec^2\theta}\end{align}\)

OpenStudy (anonymous):

thank you